# Advancing Set-Conditional Set Generation: Graph Diffusion for Fast Simulation of Reconstructed Particles

### v1

@@ Abstract
The computational intensity of detailed detector simulations poses a significant bottleneck in generating simulated data for collider experiments. This challenge inspires the continued development of fast simulation techniques based on machine learning to serve as efficient surrogate models. In our approach, a network generates a set of reconstructed objects conditioned on input particle sets. Building on the success of a slot-attention-based model, we present a new architecture utilizing diffusion, showcasing an enhanced performance in the context of single jets.
[…]
fastsim.bib
  • January 2024
Keywords: fast simulation, transformer, graph networks, slot-attention, conditional generation, diffusion
1 Introduction
@@
For experiments at the Large Hadron Collider (LHC), the demand for simulated data has increased throughout recent data-taking periods and will surge by roughly a factor of ten during the High-Luminosity runs [hl-lhc]. The standard simulation pipeline used at the general-purpose experiments ATLAS and CMS is designed to closely mirror that of real data [SOFT-2010-01, Hildreth:2015kps]. Initially, proton-proton collision events are generated, followed by parton shower, hadronization, and secondary decays. The set of stable, or “truth”, particles then enter the detector volume, and their material interactions are modeled on a microscopic level using Geant4 [geant, geant2, geant3]. The resulting energy deposits in the tracker and calorimeter cells seed the formation of tracks and clusters, respectively. Subsequently, a reconstruction algorithm refines these data into objects suitable for statistical analysis. Depending on the algorithm, the reconstructed “particle flow objects” (PFOs) either represent particles directly or else consist of tracks and calorimeter clusters with kinematics adjusted to maintain energy flow at the jet level.
@@
From the perspective of computational efficiency, the main constraint in the traditional simulation pipeline comes from the highly sophisticated but expensive model of particle-detector interactions provided by Geant4. In a bid to remove this constraint, various machine learning approaches are currently being explored [detsimoverview]. These rapid surrogate models aim to drastically reduce computational time while otherwise performing similarly to the traditional simulation pipeline.
@@
The majority of research efforts in this direction have concentrated on fast calorimeter simulation [detsimoverview, detsimdiff]. In this approach, the objective is to replace the slowest link in the simulation chain while preserving the other steps, including reconstruction. Utilizing the features of an incoming particle as a conditioning factor, a generative model is employed to produce the detector response, which can then be aggregated for all particles within an event [atlasfast3]. Within the scope of the Fast Calorimeter Challenge [ds1, ds2, ds3], a wide array of methods were developed and analyzed. Various architectures, including those based on Generative Adversarial Networks (GANs) [caloshowerGAN, deeptreeGAN], normalizing flows [caloflow, caloINN] and diffusion models [CaloDiffusion, caloclouds, caloscore, calograph] were introduced. These approaches represent the calorimeter’s response through diverse formats such as images, point clouds, or graphs.
Another approach involves the replacement of the entire simulation pipeline. In studies such as [jetgpt, ganLHCevent, generateLHCevent], the final high-level objects are generated based on particles originating from the hard process. This approach is highly efficient as it encompasses all simulation steps in a single process, eliminating the need for additional processing. However, it is worth noting that this method is process-dependent, necessitating optimization and retraining of the network for each analysis. An alternative strategy entails creating a generative model for jets [epicgan, jetben, jetepic, pcdroid, MPGAN, endtoend_flow]. In this scenario, the constituents of the jet are generated conditioned on both the jet features and the particle type. Additionally, in this approach, the features of the high-level object are directly mapped to the hard-process parton. Unlike the previous method, this approach does not require adjustment for different analyses but does necessitate event-splitting and additional processing for other objects within the event.
In the pursuit of a process-independent, full-event method, a novel strategy for fast simulation known as FlashSim [flashsim] has been introduced. This method involves utilizing stable particles to predict high-level objects such as jets, fat jets, muons, electrons, and more. For each object, a separate network is trained using distinct input and target variables.
In this study, our focus lies on developing a process-independent, end-to-end, fast simulation technique that generates reconstructed particles utilizing stable particles as input. This presents a set-to-set problem. Initially, a prototype for this approach was presented in [fastsim] using a simplified version of the problem. This initial model solely considered charged particles within a jet, with smeared tracks as targets. The architecture incorporated graph neural networks with slot attention mechanisms. Now, transitioning towards a more realistic approach that encompasses both neutral and charged particles while targeting reconstructed particles, we introduce Graph Diffusion as a novel architecture, showcasing enhanced performance compared to the baseline slot attention model.
2 Dataset
@@
The network takes as input a single jet of truth particles entering the detector. For the generation task, the target is the set of reconstructed particles. Instead of relying on a parametrized smearing model (e.g. Delphes [delphes]), we obtain reconstructed particles using a realistic detector simulation followed by a particle flow algorithm described below. Each input or output reconstructed particle is represented by its momentum, direction, and charge ($p_{\text{T}}$, $\eta$, $\phi$, $|q|$). Notably, charge prediction is a new feature introduced in this study, as our previous work utilized a toy model focusing solely on charged particles. Generally, reconstructed particles are classified into five classes: charged hadrons, neutral hadrons, photons, electrons, and muons. Given the significant class imbalance within the dataset, we simplify the task by predicting only whether the object is neutral or charged.
2.1 Truth Event Generation
@@
We concentrate on a localized reconstruction of particles within a single jet. For this purpose, we utilize Pythia8 [pythia] to generate a single quark with momentum ranging between 10 GeV and 200 GeV, with the initial direction randomly selected within the ranges $|\eta|<2.5$ and $-\pi<\phi<\pi$. Subsequent to parton shower and hadronization, only stable particles with momentum exceeding 1 GeV and $|\eta|<3.0$ are retained. The set of particles contained within the jets exhibits an average cardinality of $N=6.7$, with a maximum of 25 particles and a minimum of 1 particle per truth event, as measured on the whole dataset.
2.2 Detector simulation
@@
The detector simulation uses COCOA [cocoa], a Geant4-based configurable calorimeter simulation toolkit. The detector comprises three electromagnetic and three hadronic calorimeter layers, simulating ATLAS-like materials. Geometrically, the coverage is divided into a barrel region ($|\eta|<1.5$) and two endcaps ($3.01.5<|\eta|<$ 3.0). The cell depth follows a $1/\cosh{\eta}$ profile to maintain a constant effective interaction depth across $\eta$. Moreover, the tracking region of the detector is subjected to a magnetic field of 3.8 T. To avoid particles created upstream of the calorimeter, our detector simulation is simplified by assuming that the tracker contains no material.
@@
2.3 Particle Reconstruction
To reconstruct particles, we utilize a recently developed machine learning algorithm called HGPflow [HGpflow]. The design principle of HGPflow is to learn the energy assignment between the input set of tracks and calorimeter clusters and an output set of particles using a hypergraph prediction network [hypergraph]. We trained a version of HGPflow on a statistically independent dataset of 294k single jets and evaluated this model to obtain a set of reconstructed particles for each jet in our training dataset. These predictions serve as the target for our fast simulation models.
2.4 Preprocessing
@@
3 Fast simulation algorithms
In this study, we evaluate the performance of two set-conditional set generation approaches. Both methods rely on graph neural networks, as they can naturally represent unordered, set-valued data through a graph structure. Currently, we use fully connected graphs because we have single jets. This can be easily generalized when considering full events where different event regions are loosely or not at all connected. Edges within the graph allow inter-particle relations to be encoded through message passing. Simplified illustrations of the network architectures are depicted in Fig. 1.
Figure 1: Simplified sketches of the two proposed network architectures for the set generation task. The Graph Diffusion extends the Slot Attention approach presented in detail in [fastsim] with an iterative denoising procedure.
3.1 Slot Attention (SA)
The slot-attention approach, initially proposed in [fastsim], constitutes the baseline model. In this approach, the generation task is divided into two stages: cardinality and feature prediction. Cardinality is inferred from the updated representations of truth particles following message passing. Subsequently, based on the predicted cardinality, particles are initialized with random noise and iteratively refined through a slot-attention block [slotattention], which attends to the updated truth particle representations. The introduction of noise during initialization is crucial to prevent deterministic predictions. However, in our previous and current work, we recognize that the learned noise model exhibits room for improvement, indicated by the limited precision of the predictions.
The newly introduced Graph Diffusion, detailed in the subsequent section, aims to address the limitations of the slot-attention network.
3.2 Graph Diffusion (GD)
In this approach, we maintain the cardinality prediction method as originally introduced in the slot-attention approach and solely modify the feature prediction task. After noting that a single slot-attention block lacks the expressiveness needed to transform complete noise into sufficiently accurate features, we opt for a different approach. This involves replacing the feature prediction component with a diffusion process. Initially, starting with a fully noised graph, we employ a slot attention-based network to iteratively predict a feature estimate. This estimate is used to update the noised graph until we obtain the final, denoised reconstructed particles.
@@ 3.2 Graph Diffusion (GD)
ODE solver Heun’s $2^{\text{nd}}$ order method
@@
Table 1: Parameters of the diffusion model formulation.
3.2.1 Diffusion formulation
@@
Our Graph Diffusion approach follows the EDM strategy introduced in [karras_2022] for score-based denoising diffusion models. During training, we sample noise rates $\sigma$ from the log-normal distribution, defined by $\log(\sigma)\sim\mathcal{N}(-0.8,0.8)$. The network receives scaled noisy data $x=c_{in}(\sigma)(x_{0}+\sigma\epsilon)$ as input, where $\epsilon\sim\mathcal{N}(0,1)$. To enhance network predictions, we combine the neural network output $F_{\theta}(x;\frac{1}{4}\ln{\sigma},c)$ with a skip connection $D_{\theta}(x;\sigma,c)=c_{skip}(\sigma)x+c_{out}(\sigma)F_{\theta}(x;\frac{1}{4}\ln{\sigma},c)$, where $c$ denotes contextual information. The weighted loss function minimized during training is then defined as:
$\mathcal{L}=E_{\sigma,x_{0},\epsilon,c}\big{[}\lambda(\sigma)\left\|D_{\theta}(x;\sigma,c)-x_{0}\right\|^{2}\big{]}$ (1)
@@
We adopted specific design choices from [karras_2022], as detailed in Tab. 1. During the inference process, we utilize the network predictions to compute the score function $\nabla_{x}\log p(x;\sigma,c)=(D_{\theta}(x;\sigma,c)-x)/\sigma^{2}$ and solve the diffusion ordinary differential equation (ODE) using Heun’s $2^{\text{nd}}$ order method [Heun].
@@
To improve sampling quality, we incorporate self-conditioning on the network output from the previous timestep, following a technique proposed in [chen_2023]. This involves concatenating $x_{t}$ with the previously estimated $\tilde{x}_{0}$. For training the denoising function $D_{\theta}(x,\tilde{x}_{0};\sigma,c)$ in this context, we set $\tilde{x}_{0}=0$ with a probability of $p=0.5$, reverting to the model without self-conditioning. The remaining time, we first estimate $\tilde{x}_{0}=D_{\theta}(x,0;\sigma,c)$ and then employ it for self-conditioning.
3.2.2 Architecture description
@@
Figure 2: Schematic representation of training and generation processes of the Graph Diffusion model. For the training, noise is applied to the particle flow objects. Using the updated truth information from the first network, the denoised properties are obtained through a slot-attention block. For generation, the network iteratively denoises the objects starting from noise using the ODE solver.
@@
    This segment remains unchanged from the slot attention approach, as depicted at the top of Fig. 1. The graph of truth particles characterized by $p_{T},\eta,\phi$, and $|q|$ is embedded using a multilayer-perceptron (MLP) and subsequently updated through one round of message-passing. An overall embedding $G(T)$ of the truth event $T$ is aggregated using the node-level representation. It is used to predict the categorical distribution of the output set cardinality via an MLP.
@@
    Denoising Network
    This component of the Graph Diffusion receives as input the set of Pflow particles with noised features $x$, concatenated with self-conditioning features $\tilde{x}_{0}$ and noise level $\sigma$. These features are embedded via an MLP, along with the global representation of the truth event $G(T)$, event scaling information, and the sinusoidal embedding of $\sigma$ as contextual information. The output embedding proceeds through multi-head Slot Attention layers [slotattention], where they function as queries, while the embedding of truth particles serves as keys and values. The final representation undergoes MLP layers along with the contextual information to obtain $F_{\theta}$. This is then combined with the skip connection to derive the denoised estimate $D_{\theta}$. The schematic representation of this part is shown in Fig. 2.
@@
    Since our goal is to predict the cardinality of the generated set of particles, we split the inference process into two steps. Initially, we predict the cardinality of the generated particles from the set of truth particles and initialize them with Gaussian noise. Subsequently, we employ Heun’s $2^{\text{nd}}$ order method in conjunction with our denoising network to iteratively sample the reconstructed particle features from the noise. Following generation, objects with a $\Delta R>0.5$ with respect to the truth-jet are removed to stabilize against outliers.
4 Training and Loss
We adopt a familiar loss function outlined in [fastsim]. The central element involves the application of the Hungarian algorithm [HUNGARIAN] for particle matching. For each pair of input and output particles from the two sets, we calculate the squared differences for the particle features $\log p_{\text{T}}$ and $\eta$. Due to the periodicity of $\phi$, we use the cosine loss $2(1-\cos\Delta\phi)$, which, when Taylor-expanded to second order, recovers the squared difference. Additionally, we include the Binary Cross Entropy (BCE) loss for the charge in the case of the Slot Attention model and the squared difference for Graph Diffusion.
Hungarian matching determines the pairing configuration that minimizes the sum of all pair losses per event. The total loss comprises this sum and the Cross Entropy (CE) loss for the cardinality prediction. We additionally weigh the losses for the diffusion model as defined in Eq. 1.

### v2

# Advancing Set-Conditional Set Generation: Diffusion Models for Fast Simulation of Reconstructed Particles

@@ Abstract
The computational intensity of detector simulation and event reconstruction poses a significant difficulty for data analysis in collider experiments. This challenge inspires the continued development of machine learning techniques to serve as efficient surrogate models. We propose a fast emulation approach that combines simulation and reconstruction. In other words, a neural network generates a set of reconstructed objects conditioned on input particle sets. To make this possible, we advance set-conditional set generation with diffusion models. Using a realistic, generic, and public detector simulation and reconstruction package (COCOA), we show how diffusion models can accurately model the complex spectrum of reconstructed particles inside jets.
I Introduction
@@
For experiments at the Large Hadron Collider (LHC), the demand for simulated data has increased throughout recent data-taking periods and will surge by roughly a factor of ten during the High-Luminosity runs [1]. The standard simulation pipeline used at the general-purpose experiments ATLAS and CMS is designed to closely mirror that of real data [2, 3]. Initially, proton-proton collision events are generated, followed by parton shower, hadronization, and secondary decays. The set of stable, or “truth”, particles then enter the detector volume, and their material interactions are modeled on a microscopic level using Geant4 [4, 5, 6]. The resulting energy deposits in the tracker and calorimeter cells seed the formation of tracks and clusters, respectively. Subsequently, a reconstruction algorithm refines these data into objects suitable for statistical analysis. Depending on the algorithm, the reconstructed “particle flow objects” (PFOs) either represent particles directly or else consist of tracks and calorimeter clusters with kinematics adjusted to maintain energy flow at the jet level.
@@
From the perspective of computational efficiency, the main constraint in the traditional simulation pipeline comes from the highly sophisticated, but expensive model of particle-detector interactions provided by Geant4. In a bid to remove this constraint, various machine learning approaches are currently being explored [7, 8]. These rapid surrogate models aim to drastically reduce computational time while otherwise performing similarly to the traditional simulation pipeline.
@@
The majority of research efforts in this direction have concentrated on fast calorimeter simulation [7, 8, 9, 10]. In this approach, the objective is to replace the slowest link in the simulation chain while preserving the other steps, including reconstruction. Utilizing the features of an incoming particle as a conditioning factor, a generative model is employed to produce the detector response, which can then be aggregated for all particles within an event [11]. Within the scope of the Fast Calorimeter Challenge [12, 13, 14], a wide array of methods have been developed and analyzed. Various architectures, including those based on Generative Adversarial Networks (GANs) [15, 16], normalizing flows [17, 18] and diffusion models [19, 20, 21, 22, 23] have been introduced. These approaches represent the calorimeter’s response through diverse formats such as images, point clouds, or graphs.
A complementary approach is to replace the entire simulation pipeline by combining event generation, detector simulation, and event reconstruction. While reconstruction has typically not been a computational bottleneck, this is changing with improvements in fast detector simulations and with increased combinatorics from higher instantaneous luminosities [24, 25]. While we are not aware of a model demonstrated for entire events at the level of individual PFOs, there are a number of proposals using high-level objects [26]. This approach is highly efficient as it encompasses all simulation steps in a single process, eliminating the need for additional processing. However, it is worth noting that this method is process-dependent, necessitating optimization and retraining of the network for each analysis. A related strategy entails creating a generative model for individual jets, the most complex part of entire events [27, 28, 29, 30, 31, 32, 33, 34]. In this scenario, the constituents of the jet are generated conditioned on both the jet features and the particle type. This approach is less process-dependent than learning entire events, but it still requires the model to learn about the physics of jet substructure and when moving to entire events, it necesitates event-splitting and additional processing for other objects within the event.
Our goal is to develop a fast and automated end-to-end strategy that starts at truth particles and ends with reconstructed particles. By only learning the detector response and not the full event generation, we strive to be process independent. Going all the way to PFOs addresses the computational bottleneck from reconstruction and keeps the dimensionality of the problem manageable. In the pursuit of a process-independent, full-event method, a novel strategy for fast simulation known as FlashSim [35, 36] has been introduced [^1] The model presented in Ref. [37] has similar goals, but there is no matching between the truth and reconstructed objects during training.. This method involves utilizing stable particles to predict high-level objects such as jets, fat jets, muons, electrons, and more. For each object, a separate network is trained using distinct input and target variables. Extending the end-to-end methodology to indiviudal truth and reconstructed particles requires going beyond all examples mentioned above because we need to generate a variable-length set of particles conditioned on another variable-length set of particles. A prototype for this approach was presented in Ref. [38] using a simplified version of the problem. This initial model solely considered charged particles within a jet, with smeared tracks as targets. The architecture incorporated graph neural networks with slot attention mechanisms. Now, transitioning towards a more realistic approach that encompasses both neutral and charged particles while targeting reconstructed particles, we explore diffusion models, showcasing enhanced performance compared to the baseline slot attention model. These models are studied using a detailed and realistic detector simulation and reconstruction. We will show that continuous normalizing flows [39] are the most effective approach of the methods we studied. This forms the bassis of a companion paper [40] that shows how an end-to-end fast simulation of the CMS experiment can outperform classical baselines.
This paper is organized as follows. Section II introduces the dataset, with both the detailed simulation and reconstruction steps. The neural network approaches are described in Sec. III and trained in Sec. IV. Numerical results are presented in Sec. V and the paper ends with conclusions and outlook in Sec. VI.
II Dataset
@@
The network takes as input a single jet of truth particles entering the detector [^2] We consider single jets here as the most complex objects; future work will explore entire events.. For the generation task, the target is the set of reconstructed particles. Instead of relying on a parametrized smearing model (e.g. Delphes [41]), we obtain reconstructed particles using a realistic detector simulation followed by a particle flow algorithm described below. Each input or output reconstructed particle is represented by its momentum, direction, and charge ($p_{\text{T}}$, $\eta$, $\phi$, $|q|$). Notably, charge prediction is a new feature introduced in this study, as our previous work utilized a toy model focusing solely on charged particles. Generally, reconstructed particles are classified into five classes: charged hadrons, neutral hadrons, photons, electrons, and muons. Given the significant class imbalance within the dataset, we simplify the task by predicting only whether the object is neutral or charged.
II.1 Truth Event Generation
@@
We concentrate on a localized reconstruction of particles within a single jet. For this purpose, we utilize Pythia8.3 [42] to generate a single quark with momentum ranging between 10 GeV and 200 GeV, with the initial direction randomly selected within the ranges $|\eta|<2.5$ and $-\pi<\phi<\pi$. Subsequent to the parton shower and hadronization, only stable particles with momentum exceeding 1 GeV and $|\eta|<3.0$ are retained. The set of particles contained within the jets exhibits an average cardinality of $N=6.7$, with a maximum of 25 particles and a minimum of 1 particle per truth event, as measured on the whole dataset.
II.2 Detector simulation
@@
The detector simulation uses COCOA [43], a Geant4-based configurable calorimeter simulation toolkit. The detector comprises three electromagnetic and three hadronic calorimeter layers, simulating ATLAS-like materials. Geometrically, the coverage is divided into a barrel region ($|\eta|<1.5$) and two endcaps ($3.01.5<|\eta|<$ 3.0). The cell depth follows a $1/\cosh{\eta}$ profile to maintain a constant effective interaction depth across $\eta$. Moreover, the tracking region of the detector is subjected to a magnetic field of 3.8 T. To avoid particles created upstream of the calorimeter, our detector simulation is simplified by assuming that the tracker contains no material.
@@
II.3 Particle Reconstruction
To reconstruct particles, we utilize a recently developed machine learning algorithm called HGPflow [44]. The design principle of HGPflow is to learn the energy assignment between the input set of tracks and calorimeter clusters and an output set of particles using a hypergraph prediction network [45]. We trained a version of HGPflow on a statistically independent dataset of 294k single jets and evaluated this model to obtain a set of reconstructed particles for each jet in our training dataset. These predictions serve as the target for our fast simulation models.
II.4 Preprocessing
@@
III Fast simulation algorithms
In this study, we evaluate the performance of two set-conditional set generation approaches. Both methods rely on graph neural networks, as they can naturally represent unordered, set-valued data through a graph structure. Currently, we use fully connected graphs because we have single jets. This can be readily generalized when considering full events where different event regions are loosely or not at all connected. Edges within the graph allow inter-particle relations to be encoded through message passing. Simplified illustrations of the network architectures are depicted in Fig. 1.
Figure 1: Simplified sketches of the two proposed network architectures for the set generation task. EDM and CFM extend the Slot Attention approach presented in detail in [38] with an iterative denoising procedure.
III.1 Slot Attention (SA)
The slot-attention approach, initially proposed in Ref. [38], constitutes the baseline model. In this approach, the generation task is divided into two stages: cardinality and feature prediction. Cardinality is inferred from the updated representations of truth particles following message passing. Subsequently, based on the predicted cardinality, particles are initialized with random noise and iteratively refined through a slot-attention block [46], which attends to the updated truth particle representations. The introduction of noise during initialization is crucial to prevent deterministic predictions. A WGAN [47]-like setup is used to train the model whereby a distribution-level distance is used to steer the generator. In our previous and current work, we recognize(d) that the learned noise model exhibits room for improvement, indicated by the limited precision of the predictions.
We explore diffusion models in order to address these limitations. Diffusion models are state-of-the-art in generative AI and aim to directly learn explicitly or implicity the probability density (or the gradient of the density). Details of the EDM Diffusion (EDM) and Continuous Flow Matching (CFM) approaches are detailed in the subsequent section. The CFM is based on normalizing flows, but is still a variation on diffusion models; we will refer to both EDM and CFM as ‘diffusion’ while EDM will refer specifically to the score-based approach described below.
III.2 Diffusion Models
In this approach, we maintain the cardinality prediction method as originally introduced in the slot-attention approach and solely modify the feature prediction task. After noting that a single slot-attention block lacks the expressiveness needed to transform complete noise into sufficiently accurate features, we opt for a different approach. This involves replacing the feature prediction component with a diffusion process or flow matching. Starting with a fully noised set of particles, we use a transformer-based network to iteratively predict a feature estimate in the case of diffusion or a vector field in the case of the flow matching objective. This output is used to update the noised set until we achieve the final, denoised reconstructed particles.
@@ III.2 Diffusion Models
ODE solver 4-th order PNDM [48] method
@@
Table 1: Parameters of the EDM diffusion model formulation.
III.2.1 EDM Formulation
@@
We adapt the EDM strategy introduced in Ref. [49] for score-based denoising diffusion models. During training, we sample noise rates $\sigma$ from the log-normal distribution, defined by $\log(\sigma)\sim\mathcal{N}(-0.8,0.8)$. The network receives scaled noisy data $x=c_{in}(\sigma)(x_{0}+\sigma\epsilon)$ as input, where $\epsilon\sim\mathcal{N}(0,1)$. To enhance network predictions, we combine the neural network output $F_{\theta}(x;\frac{1}{4}\ln{\sigma},c)$ with a skip connection $D_{\theta}(x;\sigma,c)=c_{skip}(\sigma)x+c_{out}(\sigma)F_{\theta}(x;\frac{1}{4}\ln{\sigma},c)$, where $c$ denotes contextual information. The weighted loss function minimized during training is then defined as:
$\mathcal{L}=E_{\sigma,x_{0},\epsilon,c}\big{[}\lambda(\sigma)\left\|D_{\theta}(x;\sigma,c)-x_{0}\right\|^{2}\big{]}\,.$ (1)
@@
We adopted specific design choices from Ref. [49], as detailed in Tab. 1. During the inference process, we utilize the network predictions to compute the score function $\nabla_{x}\log p(x;\sigma,c)=(D_{\theta}(x;\sigma,c)-x)/\sigma$ and solve the diffusion ordinary differential equation (ODE) using PNDM method from Liu et al. [48].
III.2.2 Flow Matching Formulation
Continuous normalizing flows (CNF) [39] are a type of generative model that constructs a transformation from samples $x_{0}$ of a base distribution $p_{0}$ to samples $x_{1}$ of a target distribution $p_{1}$ using an ordinary differential equation (ODE):
$\mathrm{d}x=v_{\theta}(t,x)\mathrm{d}t,$ (2)
where the vector field $v_{\theta}$ is represented by a neural network and $t$ ranges from 0 to 1. Lipman et al. [50] recently proposed a Flow Matching (FM) objective, which allows the vector field to be learned through a regression task:
$\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t,p_{t}(x)}||v_{\theta}(t,x)-u_{t}(x)||^{2}\,.$ (3)
The function $u_{t}$ defines the desired probability path between $p_{0}$ and $p_{1}$, though it is generally not known in a closed form. This issue can be addressed with the Conditional Flow Matching (CFM) objective [50, 51, 52], optimizing which is equivalent to optimizing the FM objective. In this formulation, $u_{t}$ and the probability path $p_{t}$ are built in a sample-conditional manner:
$\mathcal{L}_{\text{CFM}}(\theta)=\mathbb{E}_{t,q(z),p_{t}(x|z)}||v_{\theta}(t,x)-u_{t}(x|z)||^{2},$ (4)
where $t\in[0,1]$, $z\sim q(z)$, and $x\sim p_{t}(x|z)$. Our approach uses the modified CFM framework from [50]. By identifying the condition $z$ with the single sample $x_{1}$ (in our context, a set of PFOs), we define:
$p_{t}(x|z)$ $=\mathcal{N}(x|tx_{1},(t\sigma-t+1)^{2})\,,$ (5)
$u_{t}(x|z)$ $=\frac{x_{1}-(1-\sigma)x}{1-(1-\sigma)t}\,,$
which outlines a linear trajectory between a standard normal distribution and a Gaussian distribution centred at $x_{1}$ with a standard deviation of $\sigma$, chosen as $10^{-4}$.
Following the strategy in Ref. [53], we sample $t$ from a power-law distribution $p(t)\propto t^{\frac{1}{1+\alpha}},\ t\in[0,1]$. Setting $\alpha=0$ results in a uniform distribution, while $\alpha>0$ gives more weight to probability paths corresponding to larger $t$ values. Our experiments showed that $\alpha=2$ achieves the best results.
III.2.3 Architecture Description
@@
The EDM and CFM models share the same transformer-based architecture; network hyperparameters are shown in Table 2. To improve sampling quality, we incorporate self-conditioning on the network output from the previous timestep, following a technique proposed in Ref. [54]. This involves concatenating $x_{t}$ with the previously estimated $\tilde{x}/\tilde{v}$. For training the network in this context, we set $\tilde{x}/\tilde{v}=0$ with a probability of $p=0.5$ reverting to the model without self-conditioning. The remaining time, first estimate $\tilde{x}/\tilde{v}$ and then utilize it for the self-conditioning with detached gradients.
@@
Figure 2: Schematic representation of training and generation processes of the EDM and CFM models. For the training, noise is applied to the particle flow objects. Using the updated truth information from the first network, the denoised properties are obtained through a cross-attention block. For generation, the network iteratively denoises the objects using the ODE solver.
@@
    This segment is similar to the slot attention approach, as depicted at the top of Fig. 1. The set of truth particles characterized by $p_{T},\eta,\phi$, and $|q|$ is enriched by a sinusoidal positional encoding, embedded using a multilayer-perceptron (MLP) and subsequently updated through two transformer encoder blocks. An overall embedding $G(T)$ of the truth event $T$ is aggregated using the node-level representation. It is used to predict the categorical distribution of the output set cardinality via an MLP.
@@
    Feature prediction
    This network component takes as input the set of PFOs with noised features $x_{t}$, combined with self-conditioning features $\tilde{x}/\tilde{v}$ and the current timestep. These features are concatenated with the sinusoidal positional encoding and embedded using an MLP. The embedding process also incorporates the global representation of the truth event $G(T)$, event scaling information, and the sinusoidal embedding of the timestep as contextual information. Then, updated features are processed through three multi-head cross-attention blocks. Each block employs two cross-attention mechanisms to first update the PFOs using the truth particles and then update the truth particles using the PFOs. The final representation is then processed by an MLP together with the contextual information to produce $F_{\theta}$ or $v_{\theta}$. In the diffusion model, the output is combined with a skip connection to generate the denoised estimate $D_{\theta}$. A schematic representation of this process is shown in Fig. 2.
@@
    Since our goal is to predict the cardinality of the generated set of particles, we split the inference process into two steps. Initially, we predict the cardinality of the generated particles from the set of truth particles and initialize them with Gaussian noise. Subsequently, we employ the PNDM method [48] in conjunction with our network to iteratively sample the reconstructed particle features from the noise. Following generation, objects with a $\Delta R>0.5$ with respect to the truth-jet are removed to stabilize against outliers.
Hyperparameters
Batch size 100
Optimizer AdamW [55]
Weight decay 0.01
Learning rate $10^{-4}$
# of epochs 150
# of time steps 25
Gradient norm clip 1.0
Max output particles 51
# of transformer blocks 2
# of cross attention blocks 3
Trainable parameters 1,742,519
Table 2: Network hyperparameters of the diffusion models.
IV Training and Loss
For the Slot Attention model, we adopt a familiar loss function outlined in Ref. [38]. The central element involves the application of the Hungarian algorithm [56] for particle matching. For each pair of input and output particles from the two sets, we calculate the squared differences for the particle features $\log p_{\text{T}}$ and $\eta$. Due to the periodicity of $\phi$, we use the cosine loss $2(1-\cos\Delta\phi)$, which recovers the squared difference when Taylor expanded to the second order. Additionally, we include the Binary Cross Entropy (BCE) loss for the charge.
Hungarian matching determines the pairing configuration that minimizes the sum of all pair losses per event. The total loss comprises this sum and the Cross-Entropy (CE) loss for the cardinality prediction.
Originally, the set approach was chosen to ensure permutation invariance, as no enforced ordering has a physical motivation. However, for the diffusion, we found that due to the sampling, better performance is achieved if ordering is applied so that every particle stays on its own trajectory. Consequently, there is no need for Hungarian matching, and we use regular Mean Squared Error (MSE) loss for the diffusion models. We also weigh the EDM model’s losses as defined in Eq. 1.